Hi guys, looking for advice. I have a Threshold Stasis 2 in for repair. The amp arrived working, but absolutely cold to the touch, even after several hours of use.
I tried to adjust the bias and it is extremely sensitive. Turning the pot clockwise, even a fraction of a fraction of a millimeter, emitter voltage jumps to 500mV, 1V+, etc. Turning it the other way, same story, 50mV, 0V.
I replaced the 5K bias pots on both channels and all of the electrolytic capacitors. But the sensitive bias adjustment remains. I cannot get the amp to the proper 50C temperature. It’s either too low and cool to the touch, or jumps dangerously high. It’s so sensitive that just inserting a small screwdriver into the bias pot (without turning it) is enough to drastically change the adjustment.
The amp produces good sound and I can’t see any other problems. Any help is greatly appreciated.
I tried to adjust the bias and it is extremely sensitive. Turning the pot clockwise, even a fraction of a fraction of a millimeter, emitter voltage jumps to 500mV, 1V+, etc. Turning it the other way, same story, 50mV, 0V.
I replaced the 5K bias pots on both channels and all of the electrolytic capacitors. But the sensitive bias adjustment remains. I cannot get the amp to the proper 50C temperature. It’s either too low and cool to the touch, or jumps dangerously high. It’s so sensitive that just inserting a small screwdriver into the bias pot (without turning it) is enough to drastically change the adjustment.
The amp produces good sound and I can’t see any other problems. Any help is greatly appreciated.
Last edited:
Try changing the balance capacitor that's across the bias control BJT-A42 you might have a small amount of instability .
Update in case anyone else is having a similar issue:
Replacing the bias transistors, tantalum 0.15uF bias capacitors, and installing 21 turn bias potentiometers did not solve the extreme sensitivity of bias adjustment on either channel.
What solved the problem was replacing all wiring from the input board to the output board which is mounted to the heatsink. Each channel has 8 pieces of short wires connecting these boards, 16 pieces of wire total. Replacing all of this wiring immediately solved the problem and now bias adjusts as it should.
Replacing the bias transistors, tantalum 0.15uF bias capacitors, and installing 21 turn bias potentiometers did not solve the extreme sensitivity of bias adjustment on either channel.
What solved the problem was replacing all wiring from the input board to the output board which is mounted to the heatsink. Each channel has 8 pieces of short wires connecting these boards, 16 pieces of wire total. Replacing all of this wiring immediately solved the problem and now bias adjusts as it should.